OpenStudy (anonymous):

6√8 x 7√2.... 6√8 simplifies to 12√2, so does the answer become 19√2 or 84√2?

OpenStudy (anonymous):

You are correct in that 6sqrt(8) simplifies to 12sqrt(2). However, when you multiply that by 7sqrt2, you must multiply everything together. The square root of 2 times the square root of 2 is 2. \[6\sqrt{8}*7\sqrt{2}=6\sqrt{4}\sqrt{2}*7\sqrt{2}\] \[12\sqrt{2}*7\sqrt{2}=84*2=168\]

OpenStudy (anonymous):

OOOOH. ohk! thank you